heir pressures are 911.7 kPa, 3800 mmHg, and 117.6 psi.  


1
Expert's answer
2022-01-31T20:35:02-0500

Converting the pressure into atm

1 atm"=101.325\>kPa"

1 atm"=760mmHg"

1 atm"=14.6959\>psi"


"911.7kPa=\frac{911.7}{101.325}=8.998\>atm"


"3800mmHg=\frac{3800}{760}=5\>atm"


"117.6\>psi=\frac{117.6}{14.6959}=8.002\>atm"